OpenStudy (anonymous):

Please help! angle A in Quadrant II, tanA=-12/5 angle B in Quadrant III, sinB=-3/5 sin ( A + B) = a. 33/65 b. -33/65 c. -63/65 d. 63/65

OpenStudy (ash2326):

@IloveCharlie what's the sign of sin and cos in quadrant ii and iii ?

OpenStudy (anonymous):

I'm not sure :/ It's not part of the problem...

OpenStudy (ash2326):

ok, I'll tell you sin= +ve and cos =-ve in II quadrant sin=-ve and cos=-ve in III quadrant

OpenStudy (anonymous):

What is ve?

OpenStudy (rogue):

Well, do you know the equation for the sum of angles for sine?

OpenStudy (anonymous):

No

OpenStudy (rogue):

Sin (A + B) = sin A cos B + sin B cos A

OpenStudy (rogue):

The only thing that you can substitute into that equation at this point is that sin B = -3/5. The rest you'll have to extrapolate.

OpenStudy (rogue):

Do you know your basic trig ratios, like sine is opposite/hypotenuse, cosine is adjacent/hypotenuse, etc.?

OpenStudy (rogue):

 You know that tan A = -12/5. Tangent is opposite/adjacent. What would sine A and cosine A be then?

OpenStudy (rogue):

Ignore the negative in front of the 12, for now.

OpenStudy (anonymous):

Sine: 12/13 cosine: 5/13 ?

OpenStudy (rogue):

Yup, good =) Now is when the signs come into play. That's what ash was talking about. There is a phrase called "All Students Take Calculus" that tells you what trig function is positive/negative in what quadrant. All - all trig functions positive in Q1 Students - Sine positive in Q2, rest negative Take - Tangent positive in Q3, rest negative Calculus - Cosine positive in Q4, rest negative So based on that, can you tell me sin A and cos A again?

OpenStudy (anonymous):

... Hmm. So Q2: 12/13, -5/13 and Q3: -12/13, -5/13

OpenStudy (anonymous):

@Rogue ?

OpenStudy (rogue):

Correct for Q2, but you have to use a different triangle for angle B. You know that sin B = -3/5, so can you set up the triangle?

OpenStudy (rogue):

 So what is Cos B?